# The Panaceas for Improving Low-Rank Decomposition in Communication-Efficient Federated Learning

### v1

@@ 1 Introduction
However, FL generally requires multiple training rounds to achieve convergence, leading to considerable communication overhead. Specifically, communication occurs in two phases: the uplink, where clients transmit local model updates to the server, and the downlink, where the server sends global model parameters to the clients. Latency in both phases can significantly reduce training efficiency, especially in scenarios with limited bandwidth or large model sizes (Hönig et al., 2022). To mitigate this issue, various strategies have been proposed to compress either the uplink (Li et al., 2024a, b; Reisizadeh et al., 2020) or the bidirectional (Tang et al., 2019; Dorfman et al., 2023) communication. This paper focuses on investigating bidirectional compression methods for better training efficiency.
@@
Low-rank decomposition (Sainath et al., 2013) is an effective technique for parameter compression, which approximates a matrix by the product of smaller sub-matrices. It has been widely employed in FL for bidirectional communication compression (Yao et al., 2021; Hyeon-Woo et al., 2022). In these approaches, the server transmits a low-rank model to the clients for training and subsequently receives the optimized models from them. For example, FedHM (Yao et al., 2021) generates a low-rank model by applying truncated singular value decomposition (SVD) to the global model. However, the SVD process introduces approximation errors, causing the low-rank model to deviate from the global model. In contrast, FedLMT (Liu et al., 2024) directly trains a pre-decomposed global model, eliminating the need for SVD. Similarly, FedPara (Hyeon-Woo et al., 2022) trains a pre-decomposed model and enhances the rank of the recovered matrices using the Hadamard product. (Mei et al., 2022) further improves compression by sharing low-rank matrices across multiple layers. These methods have laid the foundation for low-rank decomposition techniques in communication-efficient FL (CEFL). Nevertheless, several challenges remain. In this paper, we address three key issues related to decomposition in FL to further enhance communication efficiency, as shown in Figure 1.
@@
Third, how to aggregate to reduce compression errors? As shown in Figure 1(c), after receiving the optimized low-rank models, the server should aggregate the recovered matrices and then obtain a new low-rank model through SVD, which explicitly introduces errors into the model parameters. Consequently, Liu et al. (2024) suggest directly aggregating the sub-matrices to avoid the SVD process. However, our analysis indicates that this aggregation introduces implicit errors. Specifically, we find that $\bar{U}(\bar{V})^{\top}\neq\bar{W}$, where $\bar{U}$, $\bar{V}$ and $\bar{W}$ are the aggregated results of the sub-matrices $U$, $V$ and the recovered matrix $W$, respectively. To address this issue, we propose Aggregation-Aware Decomposition (AAD), which decouples the multiplication of trainable sub-matrices, as shown in Figure 1(c). Through AAD, there is no discrepancy between recovering then aggregating and aggregating then recovering, i.e., $\bar{W}=\bar{U}\tilde{V}+\tilde{U}\bar{V}$, where $\tilde{U}$ and $\tilde{V}$ are fixed matrices shared among clients. AAD applies to all forms of products involving trainable parameters that require aggregation, including the Kronecker product.
@@ 3.1 Model Update Decomposition
Recent approaches in FL reduce communication overhead by training low-rank models on the client side (Hyeon-Woo et al., 2022). For example, FedHM (Yao et al., 2021) generates low-rank models by applying truncated SVD to the global model, while FedLMT (Liu et al., 2024) directly trains a pre-decomposed global model.
@@ 3.2 Block-wise Kronecker Decomposition
Decomposition is commonly applied to linear and convolutional layers. The parameter of a linear layer is a two-dimensional tensor, making low-rank decomposition straightforward. In contrast, the parameter of a convolutional layer is a four-dimensional tensor $W\in\mathbb{R}^{c_{out}\times c_{in}\times k\times k}$, which should be reshaped into a two-dimensional tensor $W^{\prime}\in\mathbb{R}^{c_{out}k\times c_{in}k}$ before decomposition. This strategy has also been utilized in (Liu et al., 2024). In BKD, we let the sub-matrices be square for simplicity, i.e., $a=b=z$, so that the restored matrix becomes $W\in\mathbb{R}^{kz^{2}\times kz^{2}}$. Here, $k^{2}$ represents the number of blocks, and $z$ is determined by $k$ and the size of the target tensor. For instance, in the case of a linear layer with dimensions $(m,n)$, $z$ is computed as $z=\lceil\sqrt[4]{mn/k^{2}}\rceil$. When $k^{2}z^{4}>mn$, only the first $mn$ parameters of $W\in\mathbb{R}^{kz^{2}\times kz^{2}}$ are reshaped into the target parameter matrix with dimensions $(m,n)$. This process is applicable to tensors of any dimension.
@@ 4 Convergence Analysis
Notations. For an $L$-layer neural network, the model parameters are represented as $\mathbf{w}=\{W_{1},\dots,W_{\rho},W_{\rho+1}+U_{\rho+1}(V_{\rho+1})^{\top},\dots,W_{L}+U_{L}(V_{L})\top\}$, where MUD is applied to the last $(L-\rho)$ layers. $R$ denotes the number of training rounds, $E$ denotes the number of iterations per round, and $T=RE$ denotes the total number of iterations. $\left\|\cdot\right\|_{2}$ denotes the $\ell_{2}$ norm of a vector and $\left\|\cdot\right\|_{F}$ denotes the Frobenius norm of a matrix. In this section, $t\in[T]$ denotes the current iteration number, rather than the training round. For the sake of analysis, we make the following assumptions, as used by FedLMT (Liu et al., 2024).
@@
At least one of the matrices $\bar{U}_{l}=\frac{1}{N}\sum_{i=1}^{N}U_{i,l}$ and $\bar{V}_{l}=\frac{1}{N}\sum_{i=1}^{N}V_{i,l}$ has a positive smallest singular value, i.e., $[\delta_{min}(\bar{U}_{l})]^{2}+[\delta_{min}(\bar{V}_{l})]^{2}\geq\psi_{uv}^{2}>0$, where $\delta_{min}(\cdot)$ denotes the smallest singular value.
Assumptions 1 and 2 are commonly used in convergence analysis under distributed settings (Li et al., 2020; Yu et al., 2019; Li et al., 2024b). Assumption 3 is also commonplace in the theoretical analysis of machine learning models (Chen et al., 2020; Liu et al., 2024). In our case, this assumption holds exactly; see Appendix C.2 for details. Given that the model dimension is finite, it is reasonable to assume that the model weights are bounded. Assumption 4 is supported by the Marchenko-Pastur theory, as discussed by Liu et al. (2024). Based on these assumptions, we present Theorem 1, which demonstrates the convergence of FedMUD with non-convex settings. The proof is provided in Appendix C.
@@
Under Assumptions 1, 2, 3 and 4, let $21<c<$ 2 be a constant and the learning rate satisfy $0<\eta\leq\min\{(\frac{\psi_{uv}^{2}}{2})^{\frac{1}{c-1}},\frac{1}{L_{s}},1\}$, with FedMUD, we have
@@ 5.1 Experimental Setup
Baselines. FedMUD is compared to several CEFL methods, including FedHM (Yao et al., 2021), FedLMT (Liu et al., 2024), FedPara (Hyeon-Woo et al., 2022), EF21-P (Gruntkowska et al., 2023) and FedBAT (Li et al., 2024b). FedHM decomposes the global model on the server through truncated SVD and sends the decomposed model to clients for local training. FedLMT directly trains a pre-decomposed global model. FedPara enhances the rank of the recovered matrix by applying the Hadamard product. EF21-P is a state-of-the-art algorithm for communication compression with the error feedback mechanism in FL. Following the original setting of EF21-P, the Rand-$K$ and Top-$K$ compressors are used to compress the local and global model updates, respectively. FedBAT is an advanced communication quantization algorithm in FL, which learns binary model updates during local training. However, it only compresses uplink communication. For a fair comparison, we also use its quantizer to compress the global model update.
Hyperparameters. The number of clients is set to 100, with 10 clients randomly selected to participate in each round of training. The local epoch is set to 3, and the batch size is 64. SGD (Bottou, 2010) is employed as the local optimizer, with the learning rate tuned from the set $\{1.0,0.3,0.1,0.03,0.01\}$. The number of training rounds is set to 100 for FMNIST and SVHN, and 200 for CIFAR-10 and CIFAR-100. The initialization of sub-matrices in the decomposition methods has a significant impact on model performance. To ensure a fair comparison, we adjust the initialization across methods to optimize performance. Specifically, the sub-matrices are randomly initialized with values drawn from the uniform distribution $U(-a,a)$, where $a$ is selected from the set $\{0.01,0.05,0.1,0.5,1,5,10\}$. In the main experiments, the compression ratio is set to $\nicefrac{{1}}{{32}}$ for all methods. For FedMUD, we also report the performance of its variants, corresponding to whether BKD or AAD is applied on top of MUD. Each experiment is run five times, and the average results are reported. The code is provided in the supplementary materials.
@@ 5.5 Ablation on Initialization Values
The initialization of sub-matrices plays a critical role in the performance of low-rank decomposition methods. In our experiments, we employed a uniform distribution $U(-a,a)$ to initialize the sub-matrices. This section examines how different initialization values affect the performance of FedMUD. In Figure 4, the $x$-axis denotes the parameter $a$, which determines the magnitude of the initialization range. As illustrated, we observe that inappropriate initialization values can lead to a decrease in model accuracy. In contrast, the proposed BKD exhibits reduced sensitivity to variations in initialization values. Further, BKD typically requires larger initialization values compared to traditional low-rank decomposition. Overall, some adjustment of the initialization values is necessary for FedMUD to achieve optimal performance. Based on these observations, we believe that it is feasible to dynamically adjust the initialization values for FedMUD during training, and we leave this as future work.
@@ 6.1 Low-Rank Decomposition
This paper investigates the application of low-rank decomposition in communication-efficient federated learning (CEFL). The work in (Konečný et al., 2016) was the first to apply low-rank decomposition for reducing communication costs, although it only compresses the uplink communication. As previously discussed, subsequent studies (Yao et al., 2021; Liu et al., 2024) face several challenges. To address these, we propose three novel techniques that significantly improve the performance of low-rank decomposition in federated learning. It is also noteworthy that low-rank decomposition has seen broad application in personalized federated learning (FL) (Hyeon-Woo et al., 2022; Liu et al., 2024; Wu et al., 2024; Dadras et al., 2024), and our proposed methods can be effectively extended to this setting to further enhance performance.
In addition, low-rank techniques have been extensively utilized for the efficient fine-tuning of large language models (LLMs). For example, Hu et al. (2022) introduced Low-Rank Adaptation (LoRA), which approximates updates to pretrained weights using products of low-rank matrices, significantly reducing memory consumption. However, the constrained rank may limit model performance. To overcome this, Lialin et al. (2024) proposed superimposing multiple low-rank matrices to increase the effective rank. Furthermore, Edalati et al. (2023) replaced matrix multiplication in LoRA with the Kronecker product, enabling theoretical reconstruction of full-rank matrices. Nevertheless, this approach is less flexible in controlling compression rates compared to the proposed BKD method.
@@ 6.2 Communication-Efficient Federated Learning
Existing methods primarily reduce communication cost in FL from two aspects: model compression and gradient compression. Model compression typically involves training a smaller model on the client side. For example, (Caldas et al., 2018; Bouacida et al., 2021) allow clients to train sub-models derived from a larger server model, while (Isik et al., 2023; Bibikar et al., 2022) train pruned models during local training. Furthermore, Yang et al. (2021) trains and communicates a binary neural network. Low-rank decomposition methods, such as FedLMT and FedHM, also fall within the model compression category. However, model compression often sacrifices the model’s expressiveness, resulting in decreased accuracy, especially when higher compression strengths are applied. In contrast, gradient compression targets model updates (i.e., accumulated gradients), minimizing the negative impact on model’s expressiveness. Existing methods typically apply pruning (Stripelis et al., 2022; Qiu et al., 2022; Li et al., 2024a) or quantization (Karimireddy et al., 2019; Li et al., 2024b) techniques to the model updates to reduce communication overhead. Notably, MUD is a form of gradient compression. By representing the model update as sub-matrices, our approach achieves significantly higher accuracy compared to existing low-rank decomposition methods. Additionally, experimental results demonstrate that our method outperforms the gradient compression methods based on pruning and quantization.
@@ Appendix A Additional Experimental Results
In this section, we present the convergence curves of different methods under various data distributions, as illustrated in Figure 6. Then, we evaluate the performance of FedMUD under the Non-IID-2 data distribution, considering various reset intervals, initialization values, and compression ratios. The experimental results, presented in Figure 7, align with the findings in Section 5. Table 4 summarizes the results of ablation experiments on decoupling and freezing, as discussed in Section 5.3, under Non-IID-2 and IID data distributions. Additionally, we provide the experimental results for all methods under the IID data distribution, as shown in Table 3. Moreover, we evaluate our method against FedAvg and FedLMT using ResNet18 (He et al., 2016) on the CIFAR-10 and TinyImageNet (Le & Yang, 2015) datasets under the Non-IID-1 setting. The compression ratio is set to 16×16$\times$16 × or 32×32$\times$32 ×, and the learning rate is tuned over $\{0.001,0.003,0.01,0.03,0.1\}$. All other experimental settings follow those described in Section 5.1. The results, summarized in Table 5, show that our method retains superior performance even with 32×32$\times$32 × communication compression, achieving accuracy closest to FedAvg.

### v2

@@ 1 Introduction
However, FL generally requires multiple training rounds to achieve convergence, leading to considerable communication overhead. Specifically, communication occurs in two phases: the uplink, where clients transmit local model updates to the server, and the downlink, where the server sends global model parameters to the clients. Latency in both phases can significantly reduce training efficiency, especially in scenarios with limited bandwidth or large model sizes (Hönig et al., 2022). To mitigate this issue, various strategies have been proposed to compress either the uplink (Li et al., 2024a, d; Reisizadeh et al., 2020) or the bidirectional (Tang et al., 2019; Dorfman et al., 2023) communication. This paper focuses on investigating bidirectional compression methods for better training efficiency.
@@
Low-rank decomposition (Sainath et al., 2013) is an effective technique for parameter compression, which approximates a matrix by the product of smaller sub-matrices. It has been widely employed in FL for bidirectional communication compression (Yao et al., 2021; Hyeon-Woo et al., 2022). In these approaches, the server transmits a low-rank model to the clients for training and subsequently receives the optimized models from them. For example, FedHM (Yao et al., 2021) generates a low-rank model by applying truncated singular value decomposition (SVD) to the global model. However, the SVD process introduces approximation errors, causing the low-rank model to deviate from the global model. In contrast, FedLMT (Liu et al., 2024a) directly trains a pre-decomposed global model, eliminating the need for SVD. Similarly, FedPara (Hyeon-Woo et al., 2022) trains a pre-decomposed model and enhances the rank of the recovered matrices using the Hadamard product. (Mei et al., 2022) further improves compression by sharing low-rank matrices across multiple layers. These methods have laid the foundation for low-rank decomposition techniques in communication-efficient FL (CEFL). Nevertheless, several challenges remain. In this paper, we address three key issues related to decomposition in FL to further enhance communication efficiency, as shown in Figure 1.
@@
Third, how to aggregate to reduce compression errors? As shown in Figure 1(c), after receiving the optimized low-rank models, the server should aggregate the recovered matrices and then obtain a new low-rank model through SVD, which explicitly introduces errors into the model parameters. Consequently, Liu et al. (2024a) suggest directly aggregating the sub-matrices to avoid the SVD process. However, our analysis indicates that this aggregation introduces implicit errors. Specifically, we find that $\bar{U}(\bar{V})^{\top}\neq\bar{W}$, where $\bar{U}$, $\bar{V}$ and $\bar{W}$ are the aggregated results of the sub-matrices $U$, $V$ and the recovered matrix $W$, respectively. To address this issue, we propose Aggregation-Aware Decomposition (AAD), which decouples the multiplication of trainable sub-matrices, as shown in Figure 1(c). Through AAD, there is no discrepancy between recovering then aggregating and aggregating then recovering, i.e., $\bar{W}=\bar{U}\tilde{V}+\tilde{U}\bar{V}$, where $\tilde{U}$ and $\tilde{V}$ are fixed matrices shared among clients. AAD applies to all forms of products involving trainable parameters that require aggregation, including the Kronecker product.
@@ 3.1 Model Update Decomposition
Recent approaches in FL reduce communication overhead by training low-rank models on the client side (Hyeon-Woo et al., 2022). For example, FedHM (Yao et al., 2021) generates low-rank models by applying truncated SVD to the global model, while FedLMT (Liu et al., 2024a) directly trains a pre-decomposed global model.
@@ 3.2 Block-wise Kronecker Decomposition
Decomposition is commonly applied to linear and convolutional layers. The parameter of a linear layer is a two-dimensional tensor, making low-rank decomposition straightforward. In contrast, the parameter of a convolutional layer is a four-dimensional tensor $W\in\mathbb{R}^{c_{out}\times c_{in}\times k\times k}$, which should be reshaped into a two-dimensional tensor $W^{\prime}\in\mathbb{R}^{c_{out}k\times c_{in}k}$ before decomposition. This strategy has also been utilized in (Liu et al., 2024a). In BKD, we let the sub-matrices be square for simplicity, i.e., $a=b=z$, so that the restored matrix becomes $W\in\mathbb{R}^{kz^{2}\times kz^{2}}$. Here, $k^{2}$ represents the number of blocks, and $z$ is determined by $k$ and the size of the target tensor. For instance, in the case of a linear layer with dimensions $(m,n)$, $z$ is computed as $z=\lceil\sqrt[4]{mn/k^{2}}\rceil$. When $k^{2}z^{4}>mn$, only the first $mn$ parameters of $W\in\mathbb{R}^{kz^{2}\times kz^{2}}$ are reshaped into the target parameter matrix with dimensions $(m,n)$. This process is applicable to tensors of any dimension.
@@ 4 Convergence Analysis
Notations. For an $L$-layer neural network, the model parameters are represented as $\mathbf{w}=\{W_{1},\dots,W_{\rho},W_{\rho+1}+U_{\rho+1}(V_{\rho+1})^{\top},\dots,W_{L}+U_{L}(V_{L})\top\}$, where MUD is applied to the last $(L-\rho)$ layers. $R$ denotes the number of training rounds, $E$ denotes the number of iterations per round, and $T=RE$ denotes the total number of iterations. $\left\|\cdot\right\|_{2}$ denotes the $\ell_{2}$ norm of a vector and $\left\|\cdot\right\|_{F}$ denotes the Frobenius norm of a matrix. In this section, $t\in[T]$ denotes the current iteration number, rather than the training round. For the sake of analysis, we make the following assumptions, as used by FedLMT (Liu et al., 2024a).
@@
At least one of the matrices $\bar{U}_{l}=\frac{1}{N}\sum_{i=1}^{N}U_{i,l}$ and $\bar{V}_{l}=\frac{1}{N}\sum_{i=1}^{N}V_{i,l}$ has a positive smallest singular value, i.e., $\delta_{min}(\bar{U}_{l})]^{2}+[\delta_{min}(\bar{V}_{l})]^{2}\geq\psi_{uv}^{2}>0$, where $\delta_{min}(\cdot)$ denotes the smallest singular value.
Assumptions 1 and 2 are commonly used in convergence analysis under distributed settings (Li et al., 2020; Yu et al., 2019; Li et al., 2024d). Assumption 3 is also commonplace in the theoretical analysis of machine learning models (Chen et al., 2020; Liu et al., 2024a). In our case, this assumption holds exactly; see Appendix D.2 for details. Given that the model dimension is finite, it is reasonable to assume that the model weights are bounded. Assumption 4 is supported by the Marchenko-Pastur theory, as discussed by Liu et al. (2024a). Based on these assumptions, we present Theorem 1, which demonstrates the convergence of FedMUD with non-convex settings. The proof is provided in Appendix D.
@@
Under Assumptions 1, 2, 3 and 4, let $1<c<2$ be a constant and the learning rate satisfy $0<\eta\leq\min\{(\frac{\psi_{uv}^{2}}{2})^{\frac{1}{c-1}},\frac{1}{L_{s}},1\}$, with FedMUD, we have
@@ 5.1 Experimental Setup
Baselines. FedMUD is compared to several CEFL methods, including FedHM (Yao et al., 2021), FedLMT (Liu et al., 2024a), FedPara (Hyeon-Woo et al., 2022), EF21-P (Gruntkowska et al., 2023) and FedBAT (Li et al., 2024d). FedHM decomposes the global model on the server through truncated SVD and sends the decomposed model to clients for local training. FedLMT directly trains a pre-decomposed global model. FedPara enhances the rank of the recovered matrix by applying the Hadamard product. EF21-P is a state-of-the-art algorithm for communication compression with the error feedback mechanism in FL. Following the original setting of EF21-P, the Rand-$K$ and Top-$K$ compressors are used to compress the local and global model updates, respectively. FedBAT is an advanced communication quantization algorithm in FL, which learns binary model updates during local training. However, it only compresses uplink communication. For a fair comparison, we also use its quantizer to compress the global model update.
Hyperparameters. The number of clients is set to 100, with 10 clients randomly selected to participate in each round of training. The local epoch is set to 3, and the batch size is 64. SGD (Bottou, 2010) is employed as the local optimizer, with the learning rate tuned from the set $\{1.0,0.3,0.1,0.03,0.01\}$. The number of training rounds is set to 100 for FMNIST and SVHN, and 200 for CIFAR-10 and CIFAR-100. The initialization of sub-matrices in the decomposition methods has a significant impact on model performance. To ensure a fair comparison, we adjust the initialization across methods to optimize performance. Specifically, the sub-matrices are randomly initialized with values drawn from the uniform distribution U(a,a)U(-a,a), where $a$ is selected from the set $\{0.01,0.05,0.1,0.5,1,5,10\}$. In the main experiments, the compression ratio is set to $\nicefrac{{1}}{{32}}$ for all methods. For FedMUD, we also report the performance of its variants, corresponding to whether BKD or AAD is applied on top of MUD. Each experiment is run five times, and the average results are reported. The code is provided in the supplementary materials.
@@ 5.5 Ablation on Initialization Values
The initialization of sub-matrices plays a critical role in the performance of low-rank decomposition methods. In our experiments, we employed a uniform distribution U(a,a)U(-a,a) to initialize the sub-matrices. This section examines how different initialization values affect the performance of FedMUD. In Figure 4, the $x$-axis denotes the parameter $a$, which determines the magnitude of the initialization range. As illustrated, we observe that inappropriate initialization values can lead to a decrease in model accuracy. In contrast, the proposed BKD exhibits reduced sensitivity to variations in initialization values. Further, BKD typically requires larger initialization values compared to traditional low-rank decomposition. Overall, some adjustment of the initialization values is necessary for FedMUD to achieve optimal performance. Based on these observations, we believe that it is feasible to dynamically adjust the initialization values for FedMUD during training, and we leave this as future work.
@@ 6.1 Low-Rank Decomposition
This paper investigates the application of low-rank decomposition in communication-efficient federated learning (CEFL). The work in (Konečný et al., 2016) was the first to apply low-rank decomposition for reducing communication costs, although it only compresses the uplink communication. As previously discussed, subsequent studies (Yao et al., 2021; Liu et al., 2024a) face several challenges. To address these, we propose three novel techniques that significantly improve the performance of low-rank decomposition in federated learning. It is also noteworthy that low-rank decomposition has seen broad application in personalized federated learning (FL) (Hyeon-Woo et al., 2022; Liu et al., 2024a; Wu et al., 2024; Dadras et al., 2024), and our proposed methods can be effectively extended to this setting to further enhance performance.
In addition, low-rank techniques have been extensively utilized for the efficient fine-tuning of large language models (LLMs). For example, Hu et al. (2022) introduced Low-Rank Adaptation (LoRA), which approximates updates to pretrained weights using products of low-rank matrices, significantly reducing memory consumption. However, the constrained rank may limit model performance. To overcome this, Lialin et al. (2024) proposed superimposing multiple low-rank matrices to increase the effective rank. Furthermore, Edalati et al. (2023) replaced matrix multiplication in LoRA with the Kronecker product, enabling theoretical reconstruction of full-rank matrices. Nevertheless, this approach is less flexible in controlling compression rates compared to the proposed BKD method. BoRA (Li et al., 2025b) employs a similar block-wise mechanism in LoRA, where block matrix multiplication with block-wise diagonal matrices is used to enhance the expressivity of LoRA. Li et al. (2025a) have also explored the initialization of low-rank matrices, and demonstrated the potential benefits of non-zero initialization.
@@ 6.2 Communication-Efficient Federated Learning
Existing methods primarily reduce communication cost in FL from two aspects: model compression and gradient compression. Model compression typically involves training a smaller model on the client side. For example, (Caldas et al., 2018; Bouacida et al., 2021) allow clients to train sub-models derived from a larger server model, while (Isik et al., 2023; Bibikar et al., 2022) train pruned models during local training. Furthermore, Yang et al. (2021) trains and communicates a binary neural network. Low-rank decomposition methods, such as FedLMT and FedHM, also fall within the model compression category. However, model compression often sacrifices the model’s expressiveness, resulting in decreased accuracy, especially when higher compression strengths are applied. In contrast, gradient compression targets model updates (i.e., accumulated gradients), minimizing the negative impact on model’s expressiveness. Existing methods typically apply pruning (Stripelis et al., 2022; Qiu et al., 2022; Li et al., 2024a) or quantization (Karimireddy et al., 2019; Li et al., 2024d) techniques to the model updates to reduce communication overhead. Notably, MUD is a form of gradient compression. By representing the model update as sub-matrices, our approach achieves significantly higher accuracy compared to existing low-rank decomposition methods. Additionally, experimental results demonstrate that our method outperforms the gradient compression methods based on pruning and quantization.
@@ Appendix A Additional Experimental Results
In this section, we present the convergence curves of different methods under various data distributions, as illustrated in Figure 6. Then, we evaluate the performance of FedMUD under the Non-IID-2 data distribution, considering various reset intervals, initialization values, and compression ratios. The experimental results, presented in Figure 7, align with the findings in Section 5. Table 4 summarizes the results of ablation experiments on decoupling and freezing, as discussed in Section 5.3, under Non-IID-2 and IID data distributions. Additionally, we provide the experimental results for all methods under the IID data distribution, as shown in Table 3. Moreover, we evaluate our method against FedAvg and FedLMT using ResNet18 (He et al., 2016) on the CIFAR-10 and TinyImageNet (Le & Yang, 2015) datasets under the Non-IID-1 setting. The compression ratio is set to $16\times$ or $32\times$, and the learning rate is tuned over $\{0.001,0.003,0.01,0.03,0.1\}$. All other experimental settings follow those described in Section 5.1. The results, summarized in Table 5, show that our method retains superior performance even with $32\times$ communication compression, achieving accuracy closest to FedAvg.